THE ELECTRIC DIPOLE MOMENT OF YTTERBIUM MONOXIDE

\underline{C. LINTON}, Physics Department, University of New Brunswick, P.O. Box 4400, Fredericton, New Brunswick, E3B 5A3, Canada; T. C. STEIMLE and D. M. GOODRIDGE, Department of Chemistry and Biochemistry, Arizona State University, Tempe, AZ 85287-1604.

Previous experimental spectroscopic data^{a} on the ytterbium monoxide molecule, YbO, has indicated that the ground state is a ^{1}Σ^{+} state originating from the closed 4f^{14} configuration of the Yb^{+2} ion. Ab initio calculations^{b} have consistently placed the low-lying states of the 4f^{13}6s superconfiguration below that of the 4f^{14} configuration. The 4f^{13}6s superconfiguration is expected to have the smaller dipole moment because of the large back-sided polarization of the 6s orbital. We therefore decided to try and measure the dipole moment in an attempt to resolve this issue and gain a better understanding of the configurational nature of the ground state. A supersonic molecular beam of ytterbium monoxide, YbO, was produced by reacting Yb atoms with oxygen in a laser ablation source, passing the beam through a pair of Stark plates and applying an electric field. The beam was interrogated by a single mode ring dye laser tuned to the R(1) line of the Ω = 0^{+} − X^{1}Σ^{+} transition at 579 nm and, from the Stark shifts of the lines, a dipole moment of ∼ 5.8 Debye was determined for both states. The details of this investigation will be presented and the results discussed in terms of the electron configurations of the two states.

aC. Linton, S. McDonald, S. Rice, M. Dulick, Y. C. Liu and R. W. Field, J. Mol. Spectrosc. 101, 332 (1983).

bM. Dolg, H. Stoll and H. Preuss, Theor. Chim. Acta, 85,441 (1993).